OpenStudy (anonymous):

please helpp Solve the equation tan 2x = 5 cot x, for 0< x < 180

OpenStudy (welshfella):

I suggest you change this to an equation in tan x using the identities tan 2x = 2 tan x / (1 - tan^2 x) and cot x = 1/ tan x

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

so we get 2 tan x = 5 ------- -- 1 - tan^ x tan x

OpenStudy (welshfella):

cross multiply 2 tan ^2 x = 5 - 5 tan^2 x now its easy

OpenStudy (welshfella):

just add 5 tan^2 x to both sides then divide both sides by 7 to give a value for tan^2 x and take the square roots